OpenStudy (christos):

A sample consists of 17 data entries and the mean is 6.5 . A new data entry, which has the value of 3, is added to the sample. Find the mean of the new sample.

OpenStudy (anonymous):

Do you want me to find the 17 data entries? /_^

OpenStudy (anonymous):

Do you need a solid 6.5? or could it come out as 6.47 and u round up the rest?

OpenStudy (anonymous):

I have 2 sets of digits where I get 6.52 and 6.47, just let me know if you want to know what they are.

OpenStudy (anonymous):

6.3055556..\[((\sum_{i=1}^{i=17}X i ) + 3)/18= ? \]

OpenStudy (anonymous):

you know \[(\sum_{1}^{17} X i )/17 = 6.5 => (\sum_{1}^{17} X i ) = 17*6.5 = 110.5 ; substitute this \in the previous equation..you will get ur new mean :D\]